# From Many Models to ONE THEORY From the Dark Ages to Enlightenment J. Iliopoulos

### v1

@@ 2 Elementary Particles: The origins
Α φεω ψεαρς εαρλιερ, ωηιλε ωορϰινγ φορ ηις τηεςις, ῞ελτμαν βεςαμε φαμιλιαρ ωιτη α μετηοδ το ςομπυτε Φεψνμαν διαγραμς υςινγ τηε ςο-ςαλλεδ “ςυττινγ ρυλες” ωηιςη ωερε προποςεδ ιν 1960 βψ Ρ.Ε. ῝υτϰοςϰψ[120]. Τηεψ ςονςιςτ ιν ςυττινγ τηε διαγραμ ιν αλλ ποςςιβλε ωαψς ανδ ρεπλαςινγ τηε ςυτ προπαγατορς βψ τηε μαςς-ςηελλ δελτα φυνςτιονς. Τηε μετηοδ γιες τηε ιμαγιναρψ παρτ οφ τηε διαγραμ ανδ ωε ςαν ρεςοερ τηε ρεαλ παρτ βψ ωριτινγ α διςπερςιον ρελατιον. ῞ελτμαν δεελοπεδ τηις αππροαςη ανδ τυρνεδ ιτ ιντο α ερψ ποωερφυλ τοολ παρτιςυλαρλψ αδαπτεδ το εξπλορε τηεοριες φορ ωηιςη τηε εντιρε φιελδ ςοντεντ ις νοτ α πριορι ϰνοων [^61] In doing so, he was concentrated in the calculation of scattering amplitudes. For Veltman all quantum field theories that give the same $S$-matrix were considered as identical. This purely semantic question was the source of endless disputes in conferences and meetings between him and many other field theorists.. Α περςοναλ εξποςιτιον οφ ηις αππροαςη ςαν βε φουνδ ιν ρεφερενςε[121]. Ηε φιρςτ ςτυδιεδ μαςςιε Ψανγ-Μιλλς τηεοριες ατ τηε ονε λοοπ λεελ[122] ανδ ςλαιμεδ ςανςελλατιονς οφ διεργενςες, αλτηουγη τηε παπερς, εςπεςιαλλψ τηε φιρςτ ονε, ωερε νοτ εαςψ το φολλοω. Ιν ρετροςπεςτ ωε ςεε τηατ ῞ελτμαν αλλοωεδ ηιμςελφ το ςηανγε τηε Φεψνμαν ρυλες ας ιφ τηε τηεορψ ωας γαυγε ιναριαντ αλτηουγη τηε μαςς τερμς βρεαϰ τηε ςψμμετρψ. Ιν φαςτ, ωιτη α ρεμαρϰαβλε ιντυιτιον, ηε ωας ρεδιςςοερινγ βψ τριαλ ανδ ερρορ τηε ονε λοοπ εφφεςτς οφ τηε Βρουτ-Ενγλερτ-Ηιγγς μεςηανιςμ. Ιν 1970, ιν ςολλαβορατιον ωιτη Η. ῞αν Δαμ, ηε ςτυδιεδ τηε ζερο μαςς λιμιτ οφ γαυγε φιελδς, βοτη φορ Ψανγ-Μιλλς ανδ γραιτψ[123] ανδ προεδ, βψ εξπλιςιτ ςαλςυλατιον, τηατ τηε λιμιτ ωας νοτ ςοντινυους βεςαυςε τηε λονγιτυδιναλ δεγρεες οφ φρεεδομ οφ τηε μαςςιε φιελδς δο νοτ δεςουπλε ωηεν τηε μαςς γοες το ζερο. Σιμιλαρ ρεςυλτς ηαδ βεεν οβταινεδ αλςο βψ Α.Α. Σλανο ανδ Λ.Δ. Φαδδεε[124], αλτηουγη τηεψ ωερε νοτ ωιδελψ ϰνοων. Ιν 1968 ῞ελτμαν τοοϰ α ςαββατιςαλ λεαε ανδ ςπεντ τηε 68-69 ψεαρ ατ Ορςαψ. Ηε λεςτυρεδ ον Ψανγ-Μιλλς φιελδς ανδ τηερε ωερε ςομε λεςτυρε νοτες ιν τηε φορμ οφ α πρεπριντ ωηιςη, ας φαρ ας Ι ϰνοω, ηας νοτ βεεν πυβλιςηεδ. Σομε οφ υς, ινςλυδινγ μψςελφ, φουνδ ιν τηεςε νοτες ουρ φιρςτ ιντροδυςτιον το Ψανγ-Μιλλς ανδ πατη ιντεγραλς.
@@
Ιν 1969 ῞ελτμαν ωας ϑοινεδ ιν ϒτρεςητ βψ Γεραρδ ´τ Ηοοφτ, α γραδυατε ςτυδεντ ωιτη ωηομ ηε ωουλδ ςηαρε τηε Νοβελ Πριζε 30 ψεαρς λατερ [^62] Initially Veltman did not want to assign to this young student the renormalisability of Yang-Mills as a thesis project because he thought it was too risky. He proposed instead the problem of $A2$ splitting which was quite popular at that time. ’t Hooft came back a bit later declaring he did not like it, so Veltman gave up and proposed Yang-Mills. In fact, ’t Hooft showed good taste because the problem of $A2$ splitting turned out to be a fake.. Ιν 1970 ´τ Ηοοφτ φολλοωεδ τηε ῝αργèςε ςυμμερ ςςηοολ – ηις αππλιςατιον το τηε Λες Ηουςηες ςυμμερ ςςηοολ ηαδ βεεν τυρνεδ δοων – ιν ωηιςη ηε ηεαρδ Β.Ω. Λεε λεςτυρινγ αβουτ τηε ρενορμαλιςατιον οφ τηε $\sigma$-μοδελ ωιτη ςποντανεους ςψμμετρψ βρεαϰινγ. Αςςορδινγ το ´τ Ηοοφτ, τηις ωας α ρεελατιον ανδ ςηεδ νεω λιγητ το τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες[127]. Ραρε αρε τηε εξαμπλες οφ ΠηΔ ςτυδεντς ωηοςε ρεςυλτς ηαδ ςυςη α προφουνδ ανδ λαςτινγ ινφλυενςε ιν πηψςιςς. Τηε ψεαρ 1971 ωας τηε αννυς μιραβιλις. ´τ Ηοοφτ πυβλιςηεδ τωο παπερς, λανδμαρϰς ιν τηε ροαδ το τηε Στανδαρδ Μοδελ[128]. Τηε φιρςτ ηας τηε τιτλε “Ρενορμαλιςατιον οφ μαςςλεςς Ψανγ-Μιλλς Φιελδς” ανδ πρεςεντς τηε φιρςτ ςομπλετε φορμυλατιον οφ τηε χυαντυμ γαυγε τηεορψ. Ιν τηις ωορϰ ´τ Ηοοφτ γενεραλιςεδ τηε Φαδδεε-Ποπο μετηοδ το αρβιτραρψ γαυγες ανδ δεριεδ τηε ςορρεςπονδινγ Φεψνμαν ρυλες. Φυρτηερμορε, ηε ινεντεδ α γαυγε ιναριαντ ρεγυλαριςατιον ςςηεμε φορ ονε-λοοπ διαγραμς βψ ιντροδυςινγ α φιςτιτιους φιφτη διμενςιον, α πρεςυρςορ οφ τηε διμενςιοναλ ρεγυλαριςατιον μετηοδ. Ωιτη τηεςε τοολς ηε γαε τηε φιρςτ ςομπλετε προοφ οφ τηε ρενορμαλιςαβιλιτψ οφ τηε τηεορψ. Τηε ςεςονδ παπερ, ωιτη τηε τιτλε “Ρενορμαλιςαβλε Λαγρανγιανς φορ Μαςςιε Ψανγ-Μιλλς Φιελδς”, ινςορπορατες τηε ΒΕΗ μεςηανιςμ ανδ ςηοως τηατ τηε ρεςυλτινγ τηεορψ ις ρενορμαλιςαβλε. Ηε υςες ηις πρειους ρεςυλτς ιν ορδερ το ςομπυτε τηε Φεψνμαν ρυλες φορ α φαμιλψ οφ γαυγες ανδ, ιν παρτιςυλαρ, τηε ονε ϰνοων ας “τηε ´τ Ηοοφτ γαυγε” ωηιςη ις μοςτ ςονενιεντ φορ ςαλςυλατιονς. Α λιμιτινγ ςαςε ιν τηις φαμιλψ ις “τηε υνιταρψ γαυγε” ιν ωηιςη τηε Λαγρανγιαν ςονταινς ονλψ τηε πηψςιςαλ δεγρεες οφ φρεεδομ. Ιτ ις νον-ρενορμαλιςαβλε βψ ποωερ ςουντινγ βυτ, βψ γαυγε ιναριανςε, ηας τηε ςαμε $S$-ματριξ ας τηε ονε ωε ςαν ςομπυτε υςινγ, φορ εξαμπλε, τηε ´τ Ηοοφτ γαυγε. Ιτ γενεραλιςες το τηε τηεορψ οφ τηε ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς τηε Στυεςϰελβεργ εφφεςτ ωε ςαω ιν ςεςτιον 1. Τηε ςομπλετε προγραμ οφ τηε Ψανγ-Μιλλς ρενορμαλιςατιον ηας βεεν πυβλιςηεδ βψ ´τ Ηοοφτ ανδ ῞ελτμαν ιν 1972[129]. Τηε φιρςτ οφ τηεςε πυβλιςατιονς ςονταινς αλςο τηε διμενςιοναλ ρεγυλαριςατιον ςςηεμε ωηιςη, ιν αδδιτιον το πρεςερε γαυγε ιναριανςε, ηας τηε αδανταγε οφ ςιμπλιφψινγ ενορμουςλψ αλλ περτυρβατιον ςαλςυλατιονς. Ιτ εξτενδς πρειους ωορϰ βψ ῝.Γ. Βολλινι ανδ Θ.Θ. Γιαμβιαγι[130].
@@
Ι ωας ιν Παρις ατ τηε ενδ οφ 1974 ωηεν Ι ρεςειεδ α τελεπηονε ςαλλ φρομ Α. Λαγαρριγυε ινιτινγ με το αν ινφορμαλ μεετινγ το διςςυςς ιμπορταντ ρεςυλτς φρομ ΣΠΕΑΡ. Γ. ´τ Ηοοφτ ωας ιςιτινγ Εςολε Νορμαλε ανδ Ι τοοϰ ηιμ αλονγ. Ιν τηοςε δαψς νεως διδ νοτ τραελ ωιτη τηε ςπεεδ οφ ιντερνετ. Β. Θεαν-Μαριε ηαδ ςομε φρομ Στανφορδ ωιτη τηε ρεςυλτς. Τηεψ ωερε ινδεεδ ερψ ιμπρεςςιε. Α 3 Γε῞ ηαδρον, δεςαψινγ ιντο πιονς, ωιτη α ωιδτη οφ λεςς τηαν 100 ϰε῞; Ινςρεδιβλε! Λαγαρριγυε αςϰεδ με ωηατ Ι τηουγητ οφ ιτ ανδ Ι ςονφεςς Ι ωας βεωιλδερεδ. Ιτ ωας ´τ Ηοοφτ ωηο φιρςτ γαε με τηε εξπλανατιον. Ιτ ις ςιμπλε το υνδερςτανδ ιν Χ῝Δ. Λετ υς ςονςιδερ τηε ςεριες οφ $1^{-}$ μεςονς: $\rho$ ις α βουνδ ςτατε οφ α χυαρϰ-αντιχυαρϰ παιρ οφ τηε φιρςτ φαμιλψ. Ιτς μαςς λαψς ωελλ αβοε τηε $2\pi$ μαςς ανδ ιτς ωιδτη ις ερψ λαργε $\sim$ 147 Με῞. Τηε $\phi$ ις αν $s\bar{s}$ βουνδ ςτατε ωιτη α μαςς οφ 1020 Με῞. Ιτ λαψς βαρελψ αβοε τηε τηρεςηολδ οφ α $K\bar{K}$ παιρ, ψετ ιτς βρανςηινγ ρατιο το $K\bar{K}$ ις 83% δεςπιτε τηε φαςτ τηατ τηε πηαςε ςπαςε ις τινψ. Τηε πυρε πιονις παρτιαλ ωιδτη ις ονλψ 650 ϰε῞. Ιν τηε ολδ δαψς ωε ηαδ α ρυλε, ςαλλεδ τηε ΟΖΙ (Οϰυβο-Ζωειγ-Ιιζυϰα) ρυλε, ονε οφ τηοςε εμπιριςαλ ρυλες οφ τηε δαρϰ αγες ωιτη νο ρεαλ τηεορετιςαλ ϑυςτιφιςατιον. Ιτ ςτατεδ τηατ ιν α χυαρϰ-αντιχυαρϰ βουνδ ςτατε, τηε δεςαψ μοδες ρεχυιρινγ τηε αννιηιλατιον οφ τηε ινιτιαλ $q\bar{q}$ παιρ, ωερε ηιγηλψ ςυππρεςςεδ [^72] It was called “the re-arrangement model” by H. Rubinstein.. Λετ υς ςομε νοω το $J/\Psi$ ανδ αςςυμε ιτ ις α $c\bar{c}$ βουνδ ςτατε. Τηε $0^{-}$ μεςονς αρε ςυπποςεδ το βε τηε πςευδο-Γολδςτονε βοςονς οφ ςποντανεουςλψ βροϰεν ςηιραλ ςψμμετρψ. Τηε λαττερ ις ερψ γοοδ φορ τηε φιρςτ φαμιλψ, χυεςτιοναβλε φορ τηε $s$ χυαρϰ, ανδ ερψ ποορ φορ ςηαρμ. Τηερεφορε ωε εξπεςτ τηε ςηαρμεδ $0^{-}$ μεςονς $D$ το βε χυιτε ηεαψ ανδ τηε μαςς οφ $J/\Psi$ το λαψ βελοω τηε $D\bar{D}$ τηρεςηολδ. Ας α ρεςυλτ $J/\Psi$ δεςαψς μαινλψ ιντο πιονς. Τηε δεςαψ αμπλιτυδε φορ α $1^{-}$ μεςον γοες τηρουγη τηρεε γλυονς, ςο τηε ωιδτη ις προπορτιοναλ το $\alpha_{s}^{3}$. Βετωεεν 1 ανδ 3 Γε῞ $\alpha_{s}$ ηας δροππεδ βψ α φαςτορ οφ τωο, ςο ωε εξπεςτ τηε $J/\Psi$ ωιδτη το βε 8 τιμες ςμαλλερ τηαν τηε $\phi$ πιονις ωιδτη. Ιτ ις πρεςιςελψ ωηατ ις φουνδ εξπεριμενταλλψ. Ας Ι ςαιδ, Ι φιρςτ ηεαρδ τηις αργυμεντ φρομ ´τ Ηοοφτ, βυτ λατερ Ι φουνδ ιτ ιν παπερς βψ Αππελχυιςτ ανδ Πολιτζερ ας ωελλ ας Δε Ρυϑυλα ανδ Γλαςηοω [170]. Τηε φιρςτ ηας α ςυβμιςςιον δατε οφ Νοεμβερ 19. Τηις μαϰες με βελιεε ιτ ωας φουνδ βεφορε τηε εξπεριμενταλ διςςοερψ.

### v2

@@ 2 Elementary Particles: The origins
Α φεω ψεαρς εαρλιερ, ωηιλε ωορϰινγ φορ ηις τηεςις, ῞ελτμαν ηαδ βεςομε φαμιλιαρ ωιτη α μετηοδ το ςομπυτε Φεψνμαν διαγραμς υςινγ τηε ςο-ςαλλεδ “ςυττινγ ρυλες” ωηιςη ωερε προποςεδ ιν 1960 βψ Ρ.Ε. ῝υτϰοςϰψ[121]. Τηεψ ςονςιςτ ιν ςυττινγ τηε διαγραμ ιν αλλ ποςςιβλε ωαψς ανδ ρεπλαςινγ τηε ςυτ προπαγατορς βψ τηε μαςς-ςηελλ δελτα φυνςτιονς. Τηε μετηοδ γιες τηε ιμαγιναρψ παρτ οφ τηε διαγραμ ανδ ωε ςαν ρεςοερ τηε ρεαλ παρτ βψ ωριτινγ α διςπερςιον ρελατιον. ῞ελτμαν δεελοπεδ τηις αππροαςη ανδ τυρνεδ ιτ ιντο α ερψ ποωερφυλ τοολ παρτιςυλαρλψ αδαπτεδ το εξπλορε τηεοριες φορ ωηιςη τηε εντιρε φιελδ ςοντεντ ις νοτ α πριορι ϰνοων [^63] In doing so, he was concentrated in the calculation of scattering amplitudes. For Veltman all quantum field theories that give the same $S$-matrix were considered as identical. This purely semantic question was the source of endless disputes in conferences and meetings between him and many other field theorists.. Α περςοναλ εξποςιτιον οφ ηις αππροαςη ςαν βε φουνδ ιν ρεφερενςε[122]. Ηε φιρςτ ςτυδιεδ μαςςιε Ψανγ-Μιλλς τηεοριες ατ τηε ονε λοοπ λεελ[123] ανδ ςλαιμεδ ςανςελλατιονς οφ διεργενςες, αλτηουγη τηε παπερς, εςπεςιαλλψ τηε φιρςτ ονε, ωερε νοτ εαςψ το φολλοω. Ιν ρετροςπεςτ ωε ςεε τηατ ῞ελτμαν αλλοωεδ ηιμςελφ το ςηανγε τηε Φεψνμαν ρυλες ας ιφ τηε τηεορψ ωας γαυγε ιναριαντ αλτηουγη τηε μαςς τερμς βρεαϰ τηε ςψμμετρψ. Ιν φαςτ, ωιτη α ρεμαρϰαβλε ιντυιτιον, ηε ωας ρεδιςςοερινγ βψ τριαλ ανδ ερρορ τηε ονε λοοπ εφφεςτς οφ τηε Βρουτ-Ενγλερτ-Ηιγγς μεςηανιςμ. Ιν 1970, ιν ςολλαβορατιον ωιτη Η. ῞αν Δαμ, ηε ςτυδιεδ τηε ζερο μαςς λιμιτ οφ γαυγε φιελδς, βοτη φορ Ψανγ-Μιλλς ανδ γραιτψ[124] ανδ προεδ, βψ εξπλιςιτ ςαλςυλατιον, τηατ τηε λιμιτ ωας νοτ ςοντινυους βεςαυςε τηε λονγιτυδιναλ δεγρεες οφ φρεεδομ οφ τηε μαςςιε φιελδς δο νοτ δεςουπλε ωηεν τηε μαςς γοες το ζερο. Σιμιλαρ ρεςυλτς ηαδ βεεν οβταινεδ αλςο βψ Α.Α. Σλανο ανδ Λ.Δ. Φαδδεε[125], αλτηουγη τηεψ ωερε νοτ ωιδελψ ϰνοων. Ιν 1968 ῞ελτμαν τοοϰ α ςαββατιςαλ λεαε ανδ ςπεντ τηε 68-69 ψεαρ ατ Ορςαψ. Ηε λεςτυρεδ ον Ψανγ-Μιλλς φιελδς ανδ τηερε ωερε ςομε λεςτυρε νοτες ιν τηε φορμ οφ α πρεπριντ ωηιςη, ας φαρ ας Ι ϰνοω, ηας νοτ βεεν πυβλιςηεδ. Σομε οφ υς, ινςλυδινγ μψςελφ, φουνδ ιν τηεςε νοτες ουρ φιρςτ ιντροδυςτιον το Ψανγ-Μιλλς ανδ πατη ιντεγραλς.
@@
Ιν 1969 ῞ελτμαν ωας ϑοινεδ ιν ϒτρεςητ βψ Γεραρδ ´τ Ηοοφτ, α γραδυατε ςτυδεντ ωιτη ωηομ ηε ωουλδ ςηαρε τηε Νοβελ Πριζε 30 ψεαρς λατερ [^64] Initially Veltman did not want to assign to this young student the renormalisability of Yang-Mills as a thesis project because he thought it was too risky. He proposed instead the problem of $A_{2}$ splitting which was quite popular at that time. ’t Hooft came back a bit later declaring he did not like it, so Veltman gave up and proposed Yang-Mills. In fact, ’t Hooft showed good taste because the problem of $A_{2}$ splitting turned out to be a fake.. Ιν 1970 ´τ Ηοοφτ φολλοωεδ τηε ῝αργèςε ςυμμερ ςςηοολ – ηις αππλιςατιον το τηε Λες Ηουςηες ςυμμερ ςςηοολ ηαδ βεεν τυρνεδ δοων – ιν ωηιςη ηε ηεαρδ Β.Ω. Λεε λεςτυρινγ αβουτ τηε ρενορμαλιςατιον οφ τηε $\sigma$-μοδελ ωιτη ςποντανεους ςψμμετρψ βρεαϰινγ. Αςςορδινγ το ´τ Ηοοφτ, τηις ωας α ρεελατιον ανδ ςηεδ νεω λιγητ το τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες[128]. Ραρε αρε τηε εξαμπλες οφ ΠηΔ ςτυδεντς ωηοςε ρεςυλτς ηαδ ςυςη α προφουνδ ανδ λαςτινγ ινφλυενςε ιν πηψςιςς. Τηε ψεαρ 1971 ωας τηε αννυς μιραβιλις. ´τ Ηοοφτ πυβλιςηεδ τωο παπερς, λανδμαρϰς ιν τηε ροαδ το τηε Στανδαρδ Μοδελ[129]. Τηε φιρςτ ηας τηε τιτλε “Ρενορμαλιςατιον οφ μαςςλεςς Ψανγ-Μιλλς Φιελδς” ανδ πρεςεντς τηε φιρςτ ςομπλετε φορμυλατιον οφ τηε χυαντυμ γαυγε τηεορψ. Ιν τηις ωορϰ ´τ Ηοοφτ γενεραλιςεδ τηε Φαδδεε-Ποπο μετηοδ το αρβιτραρψ γαυγες ανδ δεριεδ τηε ςορρεςπονδινγ Φεψνμαν ρυλες. Φυρτηερμορε, ηε ινεντεδ α γαυγε ιναριαντ ρεγυλαριςατιον ςςηεμε φορ ονε-λοοπ διαγραμς βψ ιντροδυςινγ α φιςτιτιους φιφτη διμενςιον, α πρεςυρςορ οφ τηε διμενςιοναλ ρεγυλαριςατιον μετηοδ. Ωιτη τηεςε τοολς ηε γαε τηε φιρςτ ςομπλετε προοφ οφ τηε ρενορμαλιςαβιλιτψ οφ τηε τηεορψ. Τηε ςεςονδ παπερ, ωιτη τηε τιτλε “Ρενορμαλιςαβλε Λαγρανγιανς φορ Μαςςιε Ψανγ-Μιλλς Φιελδς”, ινςορπορατες τηε ΒΕΗ μεςηανιςμ ανδ ςηοως τηατ τηε ρεςυλτινγ τηεορψ ις ρενορμαλιςαβλε. Ηε υςες ηις πρειους ρεςυλτς ιν ορδερ το ςομπυτε τηε Φεψνμαν ρυλες φορ α φαμιλψ οφ γαυγες ανδ, ιν παρτιςυλαρ, τηε ονε ϰνοων ας “τηε ´τ Ηοοφτ γαυγε” ωηιςη ις μοςτ ςονενιεντ φορ ςαλςυλατιονς. Α λιμιτινγ ςαςε ιν τηις φαμιλψ ις “τηε υνιταρψ γαυγε” ιν ωηιςη τηε Λαγρανγιαν ςονταινς ονλψ τηε πηψςιςαλ δεγρεες οφ φρεεδομ. Ιτ ις νον-ρενορμαλιςαβλε βψ ποωερ ςουντινγ βυτ, βψ γαυγε ιναριανςε, ηας τηε ςαμε $S$-ματριξ ας τηε ονε ωε ςαν ςομπυτε υςινγ, φορ εξαμπλε, τηε ´τ Ηοοφτ γαυγε. Ιτ γενεραλιςες το τηε τηεορψ οφ τηε ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς τηε Στυεςϰελβεργ εφφεςτ ωε ςαω ιν ςεςτιον 1. Τηε ςομπλετε προγραμ οφ τηε Ψανγ-Μιλλς ρενορμαλιςατιον ηας βεεν πυβλιςηεδ βψ ´τ Ηοοφτ ανδ ῞ελτμαν ιν 1972[130]. Τηε φιρςτ οφ τηεςε πυβλιςατιονς ςονταινς αλςο τηε διμενςιοναλ ρεγυλαριςατιον ςςηεμε ωηιςη, ιν αδδιτιον το πρεςερε γαυγε ιναριανςε, ηας τηε αδανταγε οφ ςιμπλιφψινγ ενορμουςλψ αλλ περτυρβατιον ςαλςυλατιονς. Ιτ εξτενδς πρειους ωορϰ βψ ῝.Γ. Βολλινι ανδ Θ.Θ. Γιαμβιαγι[131].
@@
Τηε ρεςυλτς ωερε αννουνςεδ βψ βοτη γρουπς ον Μονδαψ, Νοεμβερ 11 1974. Τηεψ αρε πυβλιςηεδ ατ τηε ςαμε ιςςυε οφ Πηψςιςαλ Ρειεω Λεττερς. Τηε Βροοϰηαεν αρτιςλε ηας ςυβμιςςιον δατε Νοεμβερ 12 ανδ τηε ΣΠΕΑΡ ονε Νοεμβερ 13. Ιν Νοεμβερ 12 Μ. Γρεςο, α τηεοριςτ φρομ Ρομε, αρριεδ ατ ΣΛΑ῝ το γιε α ςεμιναρ ανδ φουνδ τηε τηεορψ γρουπ ατ α γρεατ εξςιτεμεντ. Ηε αςϰεδ Σ. Δρελλ περμιςςιον το ςαλλ Φραςςατι ανδ γαε το Γ. Βελλεττινι τηε νεως ωιτη τηε εξαςτ ποςιτιον οφ τηε ρεςονανςε[172]. Αλτηουγη 3.1 Γε῞ ωας ατ τηε εδγε οφ ΑΔΟΝΕ´ς ενεργψ ρανγε, τηε Φραςςατι γρουπ ςουλδ ςονφιρμ τηε Αμεριςαν ρεςυλτς[173] ανδ τηειρ παπερ ωας πυβλιςηεδ ιν τηε ςαμε ιςςυε οφ ΠΡΛ [^74] It seems the G. Salvini, transmitted the paper to PRL over the phone..
Ι ωας ιν Παρις ατ τηε ενδ οφ 1974 ωηεν Ι ρεςειεδ α τελεπηονε ςαλλ φρομ Α. Λαγαρριγυε ινιτινγ με το αν ινφορμαλ μεετινγ το διςςυςς ιμπορταντ ρεςυλτς φρομ ΣΠΕΑΡ. Γ. ´τ Ηοοφτ ωας ιςιτινγ Εςολε Νορμαλε ανδ Ι τοοϰ ηιμ αλονγ. Ιν τηοςε δαψς νεως διδ νοτ τραελ ωιτη τηε ςπεεδ οφ ιντερνετ. Β. Θεαν-Μαριε ηαδ ςομε φρομ Στανφορδ ωιτη τηε ρεςυλτς. Τηεψ ωερε ινδεεδ ερψ ιμπρεςςιε. Α 3 Γε῞ ηαδρον, δεςαψινγ ιντο πιονς, ωιτη α ωιδτη οφ λεςς τηαν 100 ϰε῞; Ινςρεδιβλε! Λαγαρριγυε αςϰεδ με ωηατ Ι τηουγητ οφ ιτ ανδ Ι ςονφεςς Ι ωας βεωιλδερεδ. Ιτ ωας ´τ Ηοοφτ ωηο φιρςτ γαε με τηε εξπλανατιον. Ιτ ις ςιμπλε το υνδερςτανδ ιν Χ῝Δ. Λετ υς ςονςιδερ τηε ςεριες οφ $1^{-}$ μεςονς: $\rho$ ις α βουνδ ςτατε οφ α χυαρϰ-αντιχυαρϰ παιρ οφ τηε φιρςτ φαμιλψ. Ιτς μαςς λαψς ωελλ αβοε τηε $2\pi$ μαςς ανδ ιτς ωιδτη ις ερψ λαργε $\sim$ 147 Με῞. Τηε $\phi$ ις αν $s\bar{s}$ βουνδ ςτατε ωιτη α μαςς οφ 1020 Με῞. Ιτ λαψς βαρελψ αβοε τηε τηρεςηολδ οφ α $K\bar{K}$ παιρ, ψετ ιτς βρανςηινγ ρατιο το $K\bar{K}$ ις 83% δεςπιτε τηε φαςτ τηατ τηε πηαςε ςπαςε ις τινψ. Τηε πυρε πιονις παρτιαλ ωιδτη ις ονλψ 650 ϰε῞. Ιν τηε ολδ δαψς ωε ηαδ α ρυλε, ςαλλεδ τηε ΟΖΙ (Οϰυβο-Ζωειγ-Ιιζυϰα) ρυλε, ονε οφ τηοςε εμπιριςαλ ρυλες οφ τηε δαρϰ αγες ωιτη νο ρεαλ τηεορετιςαλ ϑυςτιφιςατιον. Ιτ ςτατεδ τηατ ιν α χυαρϰ-αντιχυαρϰ βουνδ ςτατε, τηε δεςαψ μοδες ρεχυιρινγ τηε αννιηιλατιον οφ τηε ινιτιαλ $q\bar{q}$ παιρ, ωερε ηιγηλψ ςυππρεςςεδ [^75] It was called “the re-arrangement model” by H. Rubinstein.. Λετ υς ςομε νοω το $J/\Psi$ ανδ αςςυμε ιτ ις α $c\bar{c}$ βουνδ ςτατε. Τηε $0^{-}$ μεςονς αρε ςυπποςεδ το βε τηε πςευδο-Γολδςτονε βοςονς οφ ςποντανεουςλψ βροϰεν ςηιραλ ςψμμετρψ. Τηε λαττερ ις ερψ γοοδ φορ τηε φιρςτ φαμιλψ, χυεςτιοναβλε φορ τηε $s$ χυαρϰ, ανδ ερψ ποορ φορ ςηαρμ. Τηερεφορε ωε εξπεςτ τηε ςηαρμεδ $0^{-}$ μεςονς $D$ το βε χυιτε ηεαψ ανδ τηε μαςς οφ $J/\Psi$ το λαψ βελοω τηε $D\bar{D}$ τηρεςηολδ. Ας α ρεςυλτ $J/\Psi$ δεςαψς μαινλψ ιντο πιονς. Τηε δεςαψ αμπλιτυδε φορ α $1^{-}$ μεςον γοες τηρουγη τηρεε γλυονς, ςο τηε ωιδτη ις προπορτιοναλ το $\alpha_{s}^{3}$. Βετωεεν 1 ανδ 3 Γε῞ $\alpha_{s}$ ηας δροππεδ βψ α φαςτορ οφ τωο, ςο ωε εξπεςτ τηε $J/\Psi$ ωιδτη το βε 8 τιμες ςμαλλερ τηαν τηε $\phi$ πιονις ωιδτη. Ιτ ις πρεςιςελψ ωηατ ις φουνδ εξπεριμενταλλψ. Ας Ι ςαιδ, Ι φιρςτ ηεαρδ τηις αργυμεντ φρομ ´τ Ηοοφτ, βυτ λατερ Ι φουνδ ιτ ιν παπερς βψ Αππελχυιςτ ανδ Πολιτζερ ας ωελλ ας Δε Ρυϑυλα ανδ Γλαςηοω [174]. Τηε φιρςτ ηας α ςυβμιςςιον δατε οφ Νοεμβερ 19 ανδ τηε ςεςονδ Νοεμβερ 27. Τηις μαϰες με βελιεε τηε αςψμπτοτις φρεεδομ αργυμεντ ωας φουνδ εαρλιερ.
Ας εξπεςτεδ, τηε αννουνςεμεντ οφ τηε εξπεριμενταλ διςςοερψ τριγγερεδ α λαργε νυμβερ οφ τηεορετιςαλ παπερς τρψινγ το ιντερπρετ τηε αππεαρανςε οφ τηεςε ςτρανγε ρεςονανςες ωιτη ςυςη α ναρροω ωιδτη. Μανψ προποςεδ εξοτις εξπλανατιονς βυτ, γοινγ τηρουγη τηε εαρλψ λιτερατυρε, Ι φουνδ τηρεε ωιτη τηε ςορρεςτ ιδεντιφιςατιον ας $\bar{c}c$ βουνδ ςτατες. Τηεψ αρε τηε ονες οφ ρεφερενςε [174] ας ωελλ ας ονε βψ ῝.Α. Δομινγυεζ ανδ Μ. Γρεςο[175]. Ιτ ηας α ςυβμιςςιον δατε οφ Δεςεμβερ 30 βυτ, αππαρεντλψ, ιτ ις μεντιονεδ ας πρεπριντ ιν τηε ῝ΕΡΝ πρεπριντ λιςτ ωιτη δατε Νοεμβερ 18. Ιτ υςες δυαλιτψ ςυμ ρυλες το εςτιματε τηε αλυε οφ $R$ ανδ ςονςιδερς βοτη $\Psi$ ανδ $\Psi^{\prime}$ ας $\bar{c}c$ βουνδ ςτατες [^76] However, in this paper the narrow widths are erroneously attributed to weak interactions..